OpenStudy (lauluisa):

A group of friends are going to the county fair. Each ticket costs $7. Write an equation to model the total cost of the group's tickets. What would be the cost for 13 tickets? a) y=7x; $91 b) y= x+7; $20 c) y=x+7; $91 d)y=7x; $21

OpenStudy (anonymous):

A

OpenStudy (math&ing001):

Say x is the number of tickets, and y the total cost would be : y = 7*x = 7*13

OpenStudy (anonymous):

7*13=91 So it would be y=7x